The inreversible isomerization


A->B was carried out in a batch reactor and the following concentration-time data were obtained: T(min) 0 3 5 8 10 12 15 17.5


CA(mol/dm3) 4.0 2.89 2.25 1.45 1.0 0.65 0.25 0.07


(a) Use regression to determine the reaction order, a, and the specific reaction rate, ka.

1
Expert's answer
2021-12-14T03:10:01-0500

Momentum after release of steam P.F.f 

Pf = mv + m'v'       (1)

Where mass of empty launch steam m=7500 kg

m' is mass of satellite , m'=250 kg

v is final velocity of launch steam after releasing satellite, v=900 m/s

v' is final velocity of the satellite after release

Putting all the values in equation (1), we get

Pf =(7500 kg)(900 m/s) + (250 kg)v'

Pf = 6.750×10kg•m/s + (250 kg)v'


Now applying conservation of momentum

Momentum before release of satellite = momentum after release of satellite

Pi = Pf

7.750×106 kg•m/s = 6.750×10kg•m/s + (250 kg)v'

(7.750×106 - 6.750×10)kg•m/s= (250 kg)v'

106 m/s = 250v'

v' = (106/250) m/s